# Multigraph Message Passing with Bi-Directional Multi-Edge Aggregations

### v1

@@ 3.4 Bi-directional Multigraph Message Passing
Bi-directional message passing (Egressy et al. (2024)) has been shown to significantly improve model expressivity, especially for directed graphs. In standard message passing, a node does not receive information from its outgoing neighbors unless those neighbors are also incoming. Consequently, the node is unable to account for its outgoing edges.
We ensure that our proposed model, MEGA-GNN, also supports bi-directional message passing for multigraphs. In MEGA-GNN, a node is able to receive messages from its outgoing neighbors by using the reversed versions of their outgoing edges. Formally, the reversed edge set (illustrated in Figure 4(c)) is defined as ^=$\hat{{\mathcal{E}}}=\{\!\{\hat{e}_{ji}=(j,i)\ |\ (i,j)\in{\mathcal{E}}\}\!\}$, and the edge features for each reverse edge are initialized using the corresponding edge features from the original graph $\hat{\mathbf{e}}_{jip}^{(0)}=\mathbf{e}_{ijp}^{(0)}$, where $p\in\{1..P_{ij}\}$. In the reversed multigraph, artificial nodes $\hat{{\mathcal{V}}}^{art}$ are placed between each unique pair of nodes, in a manner that mirrors their positioning in the original multigraph:
@@
The incoming neighbors of the artificial node $\hat{v}_{ij}^{art}$ is given by the multiset $ME_{ji}$, the cardinality of which is $P_{ji}$. Let $\hat{\mathbf{h}}_{ij}\in\mathbb{R}^{d}$ be the latent features of the artificial node $\hat{v}_{ij}^{art}$.
$\begin{split}\mathbf{\hat{h}}_{ij}^{(l-1)}=\text{EdgeAgg}\{\mathbf{\hat{e}}_{jip}^{(l-1)}\ |\ p\in\{1..P_{ji}\}\}.\end{split}$ (8)
$\hat{\mathbf{h}}_{ij}^{(l)}$ is the aggregation of the incoming edges from node $i$ to node $j$ in the reversed multigraph,
$\begin{split}\mathbf{\hat{a}}_{j}^{(l-1)}&=\text{AGG}\{\hat{f}^{(l-1)}(\mathbf{x}_{i}^{(l-1)},\mathbf{\hat{h}}_{ij}^{(l-1)})\ |\ i\in N_{out}(j)\}\\ \mathbf{x}_{j}^{(l)}&=g_{v}^{(l-1)}\bigg{(}\mathbf{x}_{j}^{(l-1)},\mathbf{a}_{j}^{(l-1)},\mathbf{\hat{a}}_{j}^{(l-1)}\bigg{)},\\ \end{split}$ (9)
$N_{out}(j)=\{i\in{\mathcal{V}}|(j,i)\in{\mathcal{E}}^{supp}\}$ and $\hat{f}^{(l)}$ is a function that constructs the messages for the reversed edges. As in 2, the latent features of each individual edge can be updated without loss of generality:
$\begin{split}\mathbf{\hat{e}}_{ijp}^{(l)}=\hat{g}_{e}^{(l-1)}\big{(}\mathbf{x}_{i}^{(l-1)},\mathbf{\hat{h}}_{ij}^{(l-1)},\mathbf{x}_{j}^{(l-1)}\big{)},\\ \end{split}$ (10)
where $\hat{g}_{e}^{(l)}$ is the update function for the reversed edges. To compute $a_{j}^{(l-1)}$ we use the same equations given in the previous subsection.
@@ 3.5 Properties
In this section, we establish several key theoretical properties of MEGA-GNN and Multi-GNN, analyzing and comparing their expressiveness and permutation invariance under different conditions. In particular, we show that MEGA-GNN is always permutation equivariant, a property not shared by Multi-GNN. Next, we prove that MEGA-GNN is universal given a strict total order on the edges.
@@
(Permutation Equivariance). Given permutation-invariant EdgeAgg and AGG functions, the framework defined by Equations 4, 5, and 6 is permutation-equivariant.
The proof of Theorem 3.1 is provided in Appendix A.2. For neural networks operating on graph-structured data, permutation equivariance is a crucial property, ensuring that the model’s output remains consistent regardless of the ordering of nodes and edges. Our model is explicitly designed to maintain permutation equivariance at both the node and edge levels. Permutation invariance for graph-level outputs can be achieved by using a permutation-invariant graph readout function.
@@
The message-passing framework defined by Equations 4, 5, and 6 can compute unique node IDs in connected multigraphs given a strict total ordering of the edges.
@@
Universality. Given sufficient layers with sufficient expressiveness and width as well as a strict total ordering of the edges, the MEGA-GNN is universal.
@@
We emphasize that the ability to assign unique node IDs does not violate the permutation equivariance of MEGA-GNN. In practice, the model does not explicitly assign IDs or rely on node/edge indices or their order. Instead, edge features such as timestamps provide a natural strict total ordering. As shown theoretically, if such an ordering exists, MEGA-GNN can assign unique node IDs, ensuring that permutation equivariance and universality can coexist without conflict.
@@ 4 Experimental Results
This section presents our experiments designed to evaluate the performance of MEGA-GNN. We focus on financial transaction graphs, which are multigraphs, and address two key applications in financial crime analysis: illicit transaction detection through edge classification and phishing account detection in Ethereum transactions through node classification. We provide detailed descriptions of the datasets used in our experiments and the baselines against which we compare our method.
@@
Anti-Money Laundering (AML): Given the concerns over privacy and legal regulations, there is a scarcity of publicly available financial transaction data. Recently, Altman et al. (2023) release realistic synthetic financial transaction data for money laundering detection. In this synthetic data, a multi-agent world simulates transactions between entities such as banks, individuals, and companies. We use Small_LI, Small_HI, Medium_LI, and Medium_HI datasets, where LI indicates low-illicit and HI indicates high-illicit, referring to the ratios of illicit transactions in the datasets. The task is to perform binary classification on the transactions, labeling them as either illicit or non-illicit.
@@
Ethereum Phishing Transaction Network: Since access to real financial transaction data from banks is limited due to privacy concerns, cryptocurrencies provide an alternative data source. In our study, we use the Ethereum transaction dataset (Chen et al. (2021)), where accounts are treated as nodes and transactions as edges. Each node has a label indicating whether it is a phishing node or not. The edges contain attributes such as transaction amount and timestamp.
@@
Implementation: We implement our solution using PyTorch Geometric Fey & Lenssen (2019). We use the same temporal split as the Multi-GNN baseline across the AML datasets and apply a temporal split of 65-15-20 on the ETH dataset across the edges (transactions). We adopt Ego-IDs (You et al. (2021)) in our model in the edge classification experiments (see Table 2) and incorporate bi-directional message passing (see Section 3.4 for details) in both edge and node classification experiments (see Table 2 and 4). To ensure the statistical significance of our results, each experiment is repeated five times with different random seeds, and the mean and the standard deviation of these runs are reported. Further details about the models and the hyperparameters are provided in Appendix B.2 and Appendix B.1, respectively.
@@
Baselines We select several baselines for comparison. In particular, we include GFP+LightGBM and GFP+XGBoost Blanuša et al. (2024), as well as Multi-GNN Egressy et al. (2024), as the state-of-the-art approaches. We should note that the baseline method Multi-GNN also incorporates Ego-IDs (You et al. (2021)). Additionally, within the general multigraph message passing framework (see Section 3.3), we explore different aggregation methods such as GIN (Xu et al. (2019)), PNA (Corso et al. (2020)) and GenAgg (Kortvelesy et al. (2023)), can be applied to both multi-edge aggregation (see Appendix B.2 for details) and node-level aggregation. Note that our models can also use a combination of different multi-edge and node aggregation functions.
@@ 4.1 Edge Classification
We evaluate our proposed MEGA-GNN models across four different synthetic AML datasets, as shown in Table 2. Six distinct MEGA-GNN models are employed, using different combinations of aggregation functions across the two stages. For clarity, we adopt the following naming convention: MEGA-GIN refers to the model that applies GIN Xu et al. (2019) aggregation at both stages, while MEGA(GenAgg)-GIN applies GenAgg (Kortvelesy et al. (2023)) for multi-edge aggregation and GIN for node-level aggregation. The results consistently prove the effectiveness of our multi-edge aggregation mechanism in improving edge classification performance across datasets with varying sizes and levels of class imbalance, regardless of the aggregation function combinations used.
Table 2 shows that our MEGA-GNN models consistently outperform both traditional GNN baselines, such as GIN and PNA, and state-of-the-art models, such as Multi-GNN and those based on GFP, yielding significant performance benefits across all datasets. On average, our models achieve a $9.25\%$ increase in minority-class F1 score on the HI datasets and a $13.31\%$ improvement on the LI datasets over the state-of-the-art. A clear trend of enhanced performance with the incorporation of multi-edge aggregation is visible across all these datasets. For instance, on the AML Medium HI dataset, MEGA-GIN boosts performance from $58.92\%$ (Multi-GIN) to $68.83\%$, and similarly, MEGA-PNA advances the performance from $66.48\%$ (Multi-PNA) to $78.26\%$. These gains of nearly $10\%$ are consistent across all the AML datasets, underscoring the power of multi-edge aggregation.
4.2 Node Classification
@@
ADAMM-GIN Sotiropoulos et al. (2023) 34.73 $\pm$ 15.75
ADAMM-PNA Sotiropoulos et al. (2023) 37.99 $\pm$ 5.41
ADAMM-GenAgg Sotiropoulos et al. (2023) 33.10 $\pm$ 19.32
@@
Table 4: Ablation of MEGA-GNN variants. MP stands for message passing. We highlight the best two results.
Ablation Small_LI Small_HI ETH-Kaggle
MEGA-GIN w/ Ego-IDs & Bi-directional MP 43.66 $\pm$ 0.54 70.83 $\pm$ 2.18 55.19 $\pm$ 2.33
MEGA-GIN w/ Bi-directional MP 41.67 $\pm$ 1.51 72.50 $\pm$ 3.26 57.45 $\pm$ 1.14
MEGA-GIN w/ Ego-IDs 40.79 $\pm$ 1.91 69.59 $\pm$ 1.07 42.82 $\pm$ 3.34
MEGA-GIN (Unidirectional MP) 41.45 $\pm$ 2.13 69.98 $\pm$ 2.02 43.56 $\pm$ 2.67
MEGA-PNA w/ Ego-IDs & Bi-directional MP 45.07 $\pm$ 2.26 74.01 $\pm$ 1.55 60.02 $\pm$ 5.10
MEGA-PNA w/ Bi-directional MP 44.70 $\pm$ 1.20 74.98 $\pm$ 1.59 64.84 $\pm$ 1.73
MEGA-PNA w/ Ego-IDs 45.61 $\pm$ 0.36 73.61 $\pm$ 0.55 57.62 $\pm$ 1.23
MEGA-PNA (Unidirectional MP) 43.59 $\pm$ 1.79 73.65 $\pm$ 0.36 59.13 $\pm$ 0.51
@@
Our method demonstrates consistent improvements over baseline models. For instance, MEGA-GIN improves the minority class F1 score by $6.11\%$ over the Multi-GIN model, while MEGA-GenAgg surpasses GenAgg-based GNN by $16.52\%$. The best-performing model, MEGA-PNA, delivers the highest F1 score among all models, achieving a slight improvement over Multi-PNA.
@@ 4.3 Ablations and Throughput Analysis
In Table 4, we report the results of our ablation study on the impact of EGO-IDs (You et al. (2021)) and Bi-directional Message Passing (MP) (3.4) across the Small_LI, Small_HI, and ETH-Kaggle datasets. EGO-IDs assign unique identifiers to root nodes in the sampled batch, while Bi-directional MP uses the formulation explained in Section 3.4. The unidirectional multi-edge aggregation incorporates only the two-stage aggregation mechanism described in Section 3.3.
Notably, this analysis shows that MEGA-GNN outperforms most baselines even without using bi-directional message passing and EGO-IDs. Bi-directional MP slightly improves performance for the AML datasets and significantly improves it for the ETH-Kaggle dataset, highlighting the benefit of this optimization for directed multigraphs. EGO-IDs do not yield consistent improvements for AML datasets but cause a performance degradation for the ETH-Kaggle dataset.
@@
Figure 5: Throughput (transactions/sec) and minority-class F1 score (right) comparisons between MEGA-GNN and Multi-GNN variants across AML-Small_HI, AML-Small_LI, and ETH datasets.
@@ 5 Related Work
The expressiveness of GNNs is closely linked to their universality, as more expressive models can approximate a broader class of functions. Sato et al. (2021); Loukas (2020) demonstrated that incorporating unique node identifiers can make GNNs universal, though this comes at the cost of losing permutation invariance. However, Abboud et al. (2021) showed that partially randomized node features can achieve universality while preserving permutation invariance. In a related line of work, set functions, such as those proposed in Zaheer et al. (2017), have been shown to be universal under certain input constraints, and Fuchs* & Veličković* (2023) made the connection between universal set and graph functions. While not theoretically more expressive than GIN, Corso et al. (2020) demonstrated that combining arbitrary aggregation functions leads to better empirical performance. In Kortvelesy et al. (2023), a learnable aggregation functions were proposed that improve sample efficiency compared to both PNA and DeepSet-based aggregations on multiset neighborhoods.

### v2

@@ 3.4 Bi-directional Multigraph Message Passing
Bi-directional message passing (Egressy et al. (2024)) has been shown to significantly improve model expressivity, especially for directed graphs. In standard message passing, nodes cannot receive information from their outgoing neighbors, limiting their ability to account for outgoing edges (e.g., preventing them from accessing their out-degree). Receiving messages from outgoing neighbors is distinct from using undirected edges, where incoming and outgoing edges are treated identically.
In this section, we present formulations for bi-directional message passing with two-stage aggregation in multigraphs. In MEGA-GNN, nodes receive messages from outgoing neighbors by incorporating reversed versions of their outgoing edges. Formally, the reverse edge multiset is defined as ^=$\hat{{\mathcal{E}}}=\{\!\{(j,i)\ |\ (i,j)\in{\mathcal{E}}\}\!\}$ and the features of reverse edges are initialized as $\hat{\mathbf{e}}_{ijp}^{(0)}=\mathbf{e}_{jip}^{(0)}$, where $p\in[1..P_{ji}]$. To enable two-stage aggregation for reverse edges, a new set of artificial nodes $\hat{{\mathcal{V}}}^{art}$ is introduced. These nodes are placed between each unique pair of nodes in the reverse direction, mirroring their arrangement in the original multigraph (see Figure 4(c)).
@@
The incoming neighbors of the artificial node $\hat{v}_{ij}^{art}$ are given by the multiset $ME_{ji}$, with cardinality $P_{ji}$. Let $\hat{\mathbf{h}}_{ij}\in\mathbb{R}^{d}$ denote the latent features of the artificial node $\hat{v}_{ij}^{art}$.
In the first aggregation stage, outgoing parallel edges from node $j$ to node $i$ are aggregated.
$\begin{split}\hat{\mathbf{h}}_{ij}^{(l-1)}=\text{EdgeAgg}\{\!\{\hat{\mathbf{e}}_{ijp}^{(l-1)}\ |\ p\in[1..P_{ji}]\}\!\}.\end{split}$ (8)
In the second aggregation stage, node-level aggregation is performed.
$\begin{split}\hat{\mathbf{a}}_{j}^{(l-1)}&=\text{AGG}\{\!\{\hat{f}^{(l-1)}(\mathbf{x}_{i}^{(l-1)},\hat{\mathbf{h}}_{ij}^{(l-1)})\ |\ i\in N_{out}(j)\}\!\}\\ \mathbf{x}_{j}^{(l)}&=g_{v}^{(l-1)}\bigg{(}\mathbf{x}_{j}^{(l-1)},\mathbf{a}_{j}^{(l-1)},\hat{\mathbf{a}}_{j}^{(l-1)}\bigg{)},\\ \end{split}$ (9)
where $\hat{f}^{(l-1)}$ is a function that constructs the messages for the outgoing edges and $a_{j}^{(l-1)}$ is computed using the equations in the previous subsection. Therefore, in Equation 9, the destination node $j$ is updated using messages from both its incoming and outgoing neighbors. As in Equation 6, the latent features of the reverse edges are updated:
$\begin{split}\hat{\mathbf{e}}_{ijp}^{(l)}=\hat{g}_{e}^{(l-1)}\big{(}\mathbf{x}_{i}^{(l-1)},\hat{\mathbf{e}}_{ijp}^{(l-1)},\hat{\mathbf{h}}_{ij}^{(l-1)}\big{)},\\ \end{split}$ (10)
where $\hat{g}_{e}^{(l)}$ is the update function for the reverse edges.
@@ 3.5 Properties
In this section, we establish several key theoretical properties of MEGA-GNN and Multi-GNN, analyzing and comparing their expressiveness and permutation equivariance under different conditions. In particular, we show that MEGA-GNN is always permutation equivariant, a property not shared by Multi-GNN. Next, we prove that MEGA-GNN is universal given a strict total order on the edges.
@@
(Permutation Equivariance). Given permutation invariant EdgeAgg and AGG functions, the framework defined by Equations 4, 5, and 6 is permutation equivariant.
The proof of Theorem 3.1 is provided in Appendix A.2. For neural networks operating on graph-structured data, permutation equivariance is a crucial property, ensuring that the model’s output remains consistent to the ordering of nodes and edges. Our model is explicitly designed to maintain permutation equivariance at both the node and edge levels. Permutation invariance for graph-level outputs can be achieved by using a permutation-invariant graph readout function.
@@
The message passing framework defined by Equations 4, 5, and 6 can compute unique node IDs in connected multigraphs given a strict total ordering of the edges.
@@
(Universality). Given sufficient layers with sufficient expressiveness and width as well as a strict total ordering of the edges, the MEGA-GNN is universal.
@@
We emphasize that the ability to assign unique node IDs does not violate the permutation equivariance of MEGA-GNN. In practice, the model does not explicitly assign IDs or rely on node/edge indices or their order. Instead, edge features such as timestamps provide a natural strict total ordering. As shown theoretically, if such an ordering exists, MEGA-GNN has the ability to assign unique node IDs, ensuring that permutation equivariance and universality can coexist without conflict.
@@ 4 Experimental Results
This section presents our experiments designed to evaluate the performance of MEGA-GNN. We focus on financial transaction multigraphs, and address two key applications in financial crime analysis: illicit transaction detection through edge classification and phishing account detection in Ethereum transactions through node classification. We provide detailed descriptions of the datasets used in our experiments and the baselines against which we compare our method.
@@
Anti-Money Laundering (AML): Given the concerns over privacy and legal regulations, there is a scarcity of publicly available financial transaction data. Recently, Altman et al. (2023) released realistic synthetic financial transaction data for money laundering detection. In this synthetic data, a multi-agent world simulates transactions between entities such as banks, individuals, and companies. We use Small LI, Small HI, Medium LI, and Medium HI datasets, where LI indicates low-illicit and HI indicates high-illicit, referring to the ratios of illicit transactions in the datasets. The task is to perform binary classification on the transactions, labeling them as illicit or non-illicit. The edges include four attributes related to transactions: timestamp, received amount, currency type, and payment format.
@@
Ethereum Phishing Transaction Network (ETH): Since access to real financial transaction data from banks is limited due to privacy concerns, cryptocurrencies provide an alternative data source. In our study, we use the Ethereum transaction dataset (Chen et al. (2021)), where accounts are treated as nodes and transactions as edges. Each node has a label indicating whether it is a phishing node or not. The edges include two attributes related to transactions: timestamp and received amount.
@@
Implementation [^1] Our code is publicly available on https://github.com/hcagri/MEGA-GNN: We implement our solution using PyTorch Geometric Fey & Lenssen (2019). We use the same temporal split as the Multi-GNN baseline across the AML datasets and apply a temporal split of 65-15-20 on the ETH dataset across the edges (transactions). We adopt Ego-IDs (You et al. (2021)) in our model in the edge classification experiments (see Table 2) and incorporate bi-directional message passing (see Section 3.4 for details) in both edge and node classification experiments (see Table 2 and 4). To ensure the statistical significance of our results, each experiment is repeated five times with different random seeds, and the mean and the standard deviation of these runs are reported. Further details about the models and the hyperparameters are provided in Appendix B.2 and Appendix B.1, respectively.
@@
Baselines We select several baselines for comparison. In particular, we include GFP+LightGBM and GFP+XGBoost Blanuša et al. (2024), as well as Multi-GNN Egressy et al. (2024), as the state-of-the-art approaches. We should note that the baseline method Multi-GNN also incorporates Ego-IDs (You et al. (2021)). Additionally, we use ADAMM Sotiropoulos et al. (2023) as a baseline for Table 4 on the ETH dataset. Within the proposed multigraph message passing framework (see Section 3.3), we explore different aggregation methods such as GIN (Xu et al. (2019)), PNA (Corso et al. (2020)) and GenAgg (Kortvelesy et al. (2023)). These methods can be applied to both multi-edge aggregation (see Appendix B.2 for details) and node-level aggregation.
@@ 4.1 Edge Classification
ADAMM-GIN 34.73 $\pm$ 15.75
ADAMM-PNA 37.99 $\pm$ 5.41
ADAMM-GenAgg 33.10 $\pm$ 19.32
@@
Table 4: Ablation study of MEGA-GNN variants, showing the Minority-class F1 scores (%). MP stands for message passing. We highlight the best two results.
Ablation AML Small HI AML Small LI ETH
MEGA-GIN w/ Ego-IDs & Bi-directional MP 70.83 $\pm$ 2.18 43.66 $\pm$ 0.54 55.19 $\pm$ 2.33
MEGA-GIN w/ Bi-directional MP 72.50 $\pm$ 3.26 41.67 $\pm$ 1.51 57.45 $\pm$ 1.14
MEGA-GIN w/ Ego-IDs 69.59 $\pm$ 1.07 40.79 $\pm$ 1.91 42.82 $\pm$ 3.34
MEGA-GIN (Unidirectional MP) 69.98 $\pm$ 2.02 41.45 $\pm$ 2.13 43.56 $\pm$ 2.67
MEGA-PNA w/ Ego-IDs & Bi-directional MP 74.01 $\pm$ 1.55 45.07 $\pm$ 2.26 60.02 $\pm$ 5.10
MEGA-PNA w/ Bi-directional MP 74.98 $\pm$ 1.59 44.70 $\pm$ 1.20 64.84 $\pm$ 1.73
MEGA-PNA w/ Ego-IDs 73.61 $\pm$ 0.55 45.61 $\pm$ 0.36 57.62 $\pm$ 1.23
MEGA-PNA (Unidirectional MP) 73.65 $\pm$ 0.36 43.59 $\pm$ 1.79 59.13 $\pm$ 0.51
We evaluate our proposed MEGA-GNN models across four different synthetic AML datasets, as shown in Table 2. Six distinct MEGA-GNN models are employed, using different combinations of aggregation functions across the two stages. For clarity, we adopt the following naming convention: MEGA-GIN refers to the model that applies GIN Xu et al. (2019) aggregation at both stages, while MEGA(GenAgg)-GIN applies GenAgg (Kortvelesy et al. (2023)) for multi-edge aggregation and GIN for node-level aggregation.
Table 2 shows that our MEGA-GNN models consistently outperform both traditional GNN baselines, such as GIN and PNA, as well as state-of-the-art models, including Multi-GNN and GFP-based approaches, yielding significant performance improvements across all datasets. On average, our models achieve a $9.25\%$ increase in minority-class F1 score on the HI datasets and a $13.31\%$ improvement on the LI datasets compared to the state-of-the-art. A clear trend of enhanced performance with the incorporation of multi-edge aggregation is visible across all these datasets, regardless of the specific combinations of aggregation functions used. For instance, on the AML Medium HI dataset, MEGA-GIN boosts performance from $58.92\%$ (Multi-GIN) to $68.83\%$, and similarly, MEGA-PNA advances the performance from $66.48\%$ (Multi-PNA) to $78.26\%$. These gains of nearly $10\%$ are consistent across all the AML datasets, underscoring the power of multi-edge aggregation.
4.2 Node Classification
@@
Our method demonstrates consistent improvements over baseline models. For instance, MEGA-GIN improves the minority class F1 score by $6.11\%$ over the Multi-GIN model, while MEGA-GenAgg surpasses Multi-GenAgg by $16.52\%$. The best-performing model, MEGA-PNA, delivers the highest F1 score among all models, achieving a slight improvement over the state-of-the-art Multi-PNA.
@@ 4.3 Ablations and Throughput Analysis
In Table 4, we report the results of our ablation study on the impact of EGO-IDs (You et al. (2021)) and Bi-directional Message Passing across the AML Small LI, AML Small HI, and ETH datasets. EGO-IDs assign unique identifiers to root nodes in the sampled batch, while Bi-directional MP uses the formulation explained in Section 3.4. The unidirectional multi-edge aggregation applies only the two-stage aggregation mechanism (see Section 3.3) on top of the base GNN model.
Notably, this analysis shows that MEGA-GNN outperforms most baselines even without using bi-directional message passing and EGO-IDs. Bi-directional MP slightly improves performance for the AML datasets and significantly improves it for the ETH dataset, highlighting the benefit of this optimization for directed multigraphs. EGO-IDs do not yield consistent improvements for AML datasets but cause a performance degradation for the ETH dataset.
@@
Figure 5: Throughput (transactions/sec) and minority-class F1 score (right) comparisons between MEGA-GNN and Multi-GNN variants across AML Small HI, AML Small LI, and ETH datasets.
@@ 5 Related Work
The expressiveness of GNNs is closely linked to their universality, as more expressive models can approximate a broader class of functions. Sato et al. (2021) demonstrated that incorporating unique node identifiers can make GNNs universal, though it results in the loss of permutation invariance. However, Abboud et al. (2021); Loukas (2020) showed that partially randomized node features can achieve universality while preserving permutation invariance. In a related line of work, set functions, such as those proposed in Zaheer et al. (2017), have been shown to be universal under certain input constraints, and Fuchs* & Veličković* (2023) made the connection between universal set and graph functions. While not theoretically more expressive than GIN, Corso et al. (2020) demonstrated that combining arbitrary aggregation functions leads to better empirical performance. In Kortvelesy et al. (2023), a learnable aggregation function is proposed that improve sample efficiency compared to both PNA and DeepSet-based aggregations on multiset neighborhoods.